NONRESONANT TWO-PHOTON MASS-ANALYZED THRESHOLD IONIZATION AND ZERO KINETIC ENERGY PHOTOELECTRON SPECTROSCOPY OF KETENE

S. WANG, \underline{Y. J. SHI}, B. SIMARD, Z. J. JAKUBEK, M. BARNETT, Steacie Institute for Molecular Sciences, National Research Council of Canada, 100 Sussex Drive, Ottawa, Ontario, K1A 0R6, Canada; C-P. LIU, Department of Chemistry, National Tsing Hua University, 101, Sec 2 Kuang-Fu Road, Hsinchu, Taiwan, 30043, R. O. C.; K. MÜLLER-DETHLEFS, Department of Chemistry, The University of York, Heslington, York, YO10 5DD, UK.

The rotationally resolved nonresonant two-photon mass analyzed threshold ionization (MATI) spectra and zero kinetic energy (ZEKE) photoelectron spectra of CH_{2}CO^{+} and CD_{2}CO^{+} are presented. The MATI and ZEKE spectra of CH_{2}CO^{+} are dominated by the totally symmetric modes ν_{4} (C=C stretch) and ν_{2} (C=O stretch), with weaker excitation of ν_{3} (CH_{2} scissor), while those of CD_{2}CO^{+} are dominated by approximately equally intense ν_{2}, ν_{3}, and ν_{4}. This is consistent with the theoretical calculations of Takeshita^{a}, which indicate a strong coupling of ν_{3} and ν_{4} modes of CD_{2}CO^{+} in the ground state. In addition, weak excitations to the nontotally symmetric vibrations ν_{5} (CH_{2} wag), ν_{6} (C=C=O linear bend) and ν_{9} (C=C=O linear bend) are also observed in the MATI spectra of both isotopomers. Rotational structure is dominated by very strong ΔK_{a}=±1 manifolds with two orders of magnitude weaker ΔK_{a}=+3 features. Analysis of the ZEKE spectra provides the ionization potentials (77538.8±2 cm^{−1} for CH_{2}CO and 77533.7±2 cm^{−1} for CD_{2}CO) and rotational constants for the ground states of both ketene cations.

aK. Takeshita, J. Chem. Phys. 96, 1199, (1992)